OpenStudy (anonymous):

Solve the following system of equations. Enter the x-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=22 -2x+6y=3

OpenStudy (anonymous):

2(5x+2y)=? 2(22)=?

OpenStudy (anonymous):

1. 10x+4y=? 2. 44 Now how do i find y? @PFEH.1999

OpenStudy (anonymous):

ok I will tell you... and this one: 5(-2x+6y) = ? 5(3)=?

OpenStudy (anonymous):

1. -10x+30y=? 2. 15 Now what? :o @PFEH.1999

OpenStudy (anonymous):

so: 10x+4y = 44 -10x+30y = 15

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

got it, so now how do i solve for y? @PFEH.1999

OpenStudy (anonymous):

-10x + 10x + 4y + 30y = 15 + 44

OpenStudy (anonymous):

calculate it...

OpenStudy (anonymous):

0+34y=59

OpenStudy (anonymous):

I solved for y and got 1.735. would the answer be 2?? @PFEH.1999

OpenStudy (anonymous):

let me solve it...

OpenStudy (anonymous):

yes it's 1.735

OpenStudy (anonymous):

so if it says to solve to the nearest tenth wouldd it be 2? @PFEH.1999

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

it said that it was incorrect and that the correct answer is 3.7 :( @PFEH.1999

OpenStudy (anonymous):

really?let me solve it again please.

OpenStudy (anonymous):

yeah i got it wrong on my quiz ;o @PFEH.1999

OpenStudy (anonymous):

no i was write.it is 1.7...

OpenStudy (anonymous):

sorry right...not write :)

OpenStudy (anonymous):

thats what i thought! that is so weird!! @PFEH.1999 hehe its okay :)